OpenStudy (anonymous):

Solve for x: (x – 4)2 = 2

OpenStudy (karatechopper):

r u squaring that or multiplying?

OpenStudy (candyme):

\[(x-4)^2 =2\] is this your question?

OpenStudy (anonymous):

(x-4)2=2 2x-8=2 2x=8+2 2x=10 \[2x \div2=10\div2\] x=5

OpenStudy (anonymous):

dat is if your question is multiplying 2

OpenStudy (anonymous):

it is ^2

OpenStudy (anonymous):

then its: (x-4)(x+4)=2 x(x+4)-4(x+4)=2 \[x ^{2}+4x-4x-16=2\] \[x ^{2}-16=2\] \[x ^{2}=16+2\] \[x ^{2}=18\] now you square both sides and the square root will cancel the ^2 on x \[x=\sqrt{18}\] x=4.24

OpenStudy (anonymous):

or \[3\sqrt{2}\] its the same but the other answer was in decimals

OpenStudy (anonymous):

If that's an exponent, \((x-4)^2=2\), then \(x-4=\pm\sqrt{2}\), so \(x=4\pm\sqrt{2}\).

OpenStudy (anonymous):

no you solve the set that is (x-4)^2 as different of 2 squares,then you equate it to 2